THE VALUE OF X SATISFYING  EQUATION                             X=(2+(2-(2+X)1/2)1/2)1/2   ?

Grade:

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear rahul

 X=(2+(2-(2+X)1/2)1/2)1/2 

x≥ 0  ................1

square

x2 =(2+(2-(2+X)1/2)1/2)

or

x2 -2 =(2-(2+X)1/2)1/2

so x2 -2≥0

   so   x≥2   or x ≤-2

 from equation 1        x≥2

and RHS      2-(2+X)1/2 ≥ 0

   or     x≤2

 so only x=2 is possible

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin